# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v2

@@ 6.2 Anchor-based Trajectory Prediction
Anchor-based trajectory prediction methods assume that future trajectories are conditioned on planning objectives, such as goal destinations or potential paths. These methods use predefined spatial or trajectory “anchors” to guide the prediction process. Figure 12 illustrates the general pipeline for the coupled planning and prediction framework commonly found in the literature. The pipeline begins by encoding past trajectories and scene context into a context vector that captures essential features. This context vector is then utilized in the anchor-based decoding process to predict possible future trajectories, which can be guided by two types of anchors: goal anchors or trajectory anchors.
Goal anchor-based methods focus on predicting a set of feasible goal points that serve as anchors for trajectory generation [525]. These methods predict potential goals based on the current context and agent behavior [525, 526]. Once the goals are identified, trajectories are generated by conditioning on these anchors, often through a path-planning or motion-generation module. Examples include predicting key destinations, such as parking spots or intersections, and using goal-conditioned policies to create paths leading to these predicted goals. A significant advantage of goal anchor-based methods is their natural alignment with goal-driven tasks, such as reaching specific destinations. Trajectory anchor-based methods rely on predefined or learned trajectory templates that represent potential future paths [3]. These templates serve as coarse motion representations and are selected or predicted as bases for future trajectories. The anchors can either be directly used or refined further during prediction. Common examples include straight, curved, or turning path templates. Table 17 provides a summary of reviewed anchor-based prediction methods.
@@ 6.2.1 Goal Anchor-Based Prediction Methods
In goal anchor-based forecasting methods, plausible goal points are first predicted based on agent behavior and scene context, then used to guide trajectory generation via path-planning or motion-generation modules. Goals are often estimated based on static environmental features, including road topology, lane structures, or sidewalks. In road networks, they can be defined as lane endpoints, intersections, or turn exits, providing a structured basis for trajectory prediction.
Several SOTA prediction methods incorporate goal-directed planning principles to estimate potential destinations and generate goal-conditioned trajectories [536, 81, 529, 537]. By estimating plausible goals, these approaches significantly enhance trajectory prediction accuracy. The BiTraP model [176] utilizes a bidirectional decoder conditioned on predicted goals to reduce accumulated errors. The SGNet framework [532] introduces a stepwise goal estimator that infers and leverages goals at multiple temporal scales. Similarly, PECNet [529] focuses on predicting distant goals to improve the accuracy of long-term trajectory predictions.
@@ 6.2.2 Trajectory Anchor-Based Prediction
Trajectory anchor-based methods focus on predicting potential paths (trajectories) that a vehicle might take to reach its goals [538, 539]. TPNet model [294] generates feasible future trajectories using polynomial curves. Inspired by the sampling-based paradigm in vehicle motion planning [540, 541], PRIME combines planning-based approaches with learning-based prediction by directly searching for reachable paths and generating potential trajectory anchors. This hybrid approach handles complex agent-map interactions while satisfying environmental and kinematic constraints [511], leveraging the strengths of both model-based and learning-based methods for optimal performance in complex scenarios. Next, to enforce constraints during trajectory generation, Li et al. [542] proposed the planning-inspired hierarchical (PiH) framework, which employs a hierarchical lateral and longitudinal decomposition to generate path and goal intentions for agents.
6.3 Summary and Discussion
This section reviewed planning-conditioned trajectory prediction methods, focusing on their categorization and methodologies. These methods leverage motion planning principles to predict future trajectories while ensuring constraint compliance and contextual awareness. The reviewed approaches encompass model-based, learning-based, and hybrid techniques, along with anchor-based methods such as goal and trajectory anchors, each presenting distinct advantages and limitations in trajectory prediction.
Model-based approaches employ explicit mathematical representations of agent dynamics and environmental constraints to produce interpretable and physically feasible trajectories. Although reliable and constraint-compliant, these methods often lack adaptability to unstructured or highly dynamic environments. In contrast, learning-based methods excel in capturing complex interactions and multimodal behaviors by employing data-driven techniques, but their reliance on large datasets and inability to explicitly enforce safety constraints remain significant challenges. Anchor-based methods, including goal and trajectory anchors, offer a structured approach to trajectory prediction by aligning with real-world driving objectives. While effective in generating diverse trajectory options, these methods may encounter scalability and computational efficiency issues in dense, crowded scenarios.
@@ 6.3 Summary and Discussion
Future research in planning-conditioned trajectory prediction should address current limitations by developing hybrid frameworks that combine the robustness and interpretability of model-based methods with the adaptability of learning-based approaches, e.g., PRIME [511]. Another important direction is the anchor generation module, which typically analyzes scene context and/or integrates spatial information to infer potential destinations or paths. However, scene context analysis can be computationally expensive and should be evaluated rigorously in terms of inference time. While map-based anchor generation can be more efficient, it must still be tested for scalability as the number of agents increases, since anchors must be generated per agent. The anchors’ generation module is part of the prediction pipeline and should be evaluated online, rather than relying on pre-generated anchors.
@@ 7 Vision-Language Models for Prediction
Vision-Language Models (VLMs) are seen as promising technology that could enable vehicles and systems to deeply understand their surroundings, marking a significant advancement in the era of intelligent and explainable transportation. VLMs [543, 544, 545, 546, 547, 548, 549, 550, 551, 552] break down the barriers between textual and visual information to effectively connect multimodal data. In the context of autonomous driving, VLMs can be roughly categorized into: Multimodal-to-Text (M2T) [553, 554, 555, 556] models, which take an image and text as input and generate text as output; Multimodal-to-Vision (M2V) [557, 558] models, where the input is an image and text pair and the output is an image; and Vision-to-Text (V2T) [559, 560] models, which take an image as input and generate text as output. This section mainly focuses on M2T models in the context of perception and prediction tasks. Readers are encouraged to look into the surveys [561, 562] for further details on other VLM configurations and downstream tasks. Figure 13 shows the general flow of a VLM-based prediction model. The model processes two main inputs: visual and textual data. While the image encoder processes visual inputs, e.g., single-camera images, multi-camera feeds, or LiDAR-range image, extracting spatial and object-level features; the text encoder interprets contextual information, which can include detailed instructions, scene descriptions, specific object inquiries, navigation commands, etc. The encoded features are then passed to a cross-modal attention mechanism [563, 564, 565], which captures the relationships between visual cues and textual instructions to generate joint representations. This representation is further processed by an autonomous driving task-specific network, e.g., as for the prediction task, by a context understanding layer and decoder. The generated future trajectory considers both visual observations and semantic guidance.
@@
This section reviews VLM-based models designed for trajectory prediction and traffic scene understanding. Given the limited number of prediction models in the literature, we have included a survey of traffic scene understanding models, considering their potential integration into prediction pipelines. The section also provides a list of available language-enhanced AD datasets with brief descriptions and discusses the challenges associated with integrating VLMs into the autonomous driving stack.
@@ 7.1 Trajectory Prediction
Table 18 summarizes reviewed VLM and LLM-based models for trajectory prediction. Yang et al. [566] proposed a two-stage evaluation framework. In the first stage, a VLM model analyzes images from six camera views and a BEV visualization with labeled agents to rank the prediction difficulty. To improve the model’s response accuracy, the VLM is provided with several examples of rankings constructed based on displacement errors from the ground truth. In the second stage, the VLM’s output is used to create a subset of training data based on the ranking, resulting in a highly representative and challenging training dataset that enhances the training efficiency of a conventional prediction model. Moon et al. [567] introduced VisionTrap that leverages human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. The model comprises four main components: a per-agent state encoder that generates agent-specific context features, a visual semantic encoder that encodes images combined with road segments, a text-driven guidance block that guides learning and understanding of detailed visual semantics, and a trajectory decoder. The study includes an inference-time analysis demonstrating the model’s suitability for real-time predictions.
@@
Keysan et al. [565] expanded the previously proposed CoverNet model [69] by incorporating a text-based scene representation. The prompt accepted by the model contains structured information about the target agent and its environment, constructed from polyline and Bézier curve lane representations. Peng et al. [568] introduced an explainable lane change prediction model (LC-LLM). The study reformulates the problem of lane change prediction as a language modeling task, processing various driving scenario data as natural language prompts. To enhance transparency and prediction reliability, the model’s Chain-of-Thought (CoT) reasoning is refined to include explanatory elements in the prompts during the inference phase. Zheng et al. [569] leveraged LLMs to improve the global understanding of traffic context for motion prediction. The study transforms complex traffic scenarios and agents’ past trajectories into visual prompts (Transportation Context Maps) along with corresponding textual prompts. TC-Maps, combined with text prompts, enable the model to generate comprehensive situational insights, extracting details such as agent intentions and scenario types. The extracted information is then encoded and integrated into a conventional motion prediction pipeline.
@@ 7.2 Traffic Scene Understanding
Table 19 summarizes reviewed VLM and LLM-based models for scene understanding. Choudhary et al. [556] introduced Talk2BEV — a VLM designed to handle diverse tasks, including predicting traffic participants’ intentions, visual and spatial reasoning, and decision-making based on visual information. Romero et al. [550] argue that using VLMs to query images via natural language addresses issues of expressivity and the need for multiple single-purpose query models, but it often returns redundant and low-quality results. To improve this, they introduced Zelda, a system that adds discriminator and synonym terms to refine user queries into natural language prompts, enhancing accuracy by filtering out low-quality frames. Mo et al. [573] designed a framework for understanding driving corner cases, built upon the foundation model InternVL2-26B and consisting of four phases. For visual perception, the framework uses DINO for 2D detection and DepthAnything for depth estimation. This detection information, along with the original image, is processed by a scene-aware retrieval module to generate contextually relevant information. The Driving Prompt Optimization module then uses this information to guide the model, suppressing hallucinations and focusing on critical scene aspects, ultimately producing responses for perception and driving recommendations. Zhou et al. [574] introduced the Embodied Language Model (ELM) to enhance traditional VLMs’ scene comprehension, supporting tasks such as object localization, detailed scene description, prediction, and event memorization. ELM incorporates space-aware pre-training for robust spatial localization and time-aware token selection to effectively process long-term video data and retrieve relevant past events. Finally, Tian et al. [575] proposed the DriveVLM system for planning in complex environments, with modules for scene description, scene analysis, and hierarchical planning. The scene description module generates linguistic representations of driving conditions and passes these to the scene analysis module, which examines objects’ motion states and behaviours that may influence vehicle actions.
@@ 7.4 Summary and Discussion
Another challenge is prompt engineering. It refers to the process of crafting input prompts that guide the model toward specific behaviors or outputs, especially by aligning visual and textual inputs. Providing relevant context within the prompt significantly affects the model’s response. For example, including scene-specific cues such as “heavy traffic ahead” or “intersection with pedestrians” can influence the model’s understanding and downstream reasoning. Prompts can also include instructions on output length or format, helping ensure concise and relevant generation. In some cases, few-shot prompting is used instead, where an example of the expected output is embedded within the prompt. Additionally, techniques such as Chain-of-Thought (CoT) prompting structure the model’s reasoning into intermediate steps, enhancing both interpretability and output quality. An efficient prompt design not only improves semantic alignment but also reduces token usage, contributing to lower inference time.
Lastly, the non-deterministic nature of some VLMs [591, 592], where the same image-text input may yield different outputs across runs, presents integration challenges in safety-critical systems. Strategies such as adjusting the temperature parameter to control output variability, applying majority voting, or averaging across the top-$k$ outputs can help manage this variability. Striking the right balance between consistency and flexibility is essential: enforcing deterministic outputs enhances system stability, while preserving generative flexibility allows the model to better handle edge cases and out-of-distribution scenarios. A rigid deterministic setup may limit adaptability, especially in complex environments. Consequently, the choice between generative and deterministic use of VLMs must be carefully evaluated in the context of trajectory prediction. Although VLMs show promise for enhancing situational understanding and predictive reasoning, their optimal configuration within the autonomous driving stack remains an on-going research investigation.
@@ 8 Collaboration between AVs for Prediction
There are certain limitations that cannot be resolved through multimodal inputs or advanced mechanisms. Each sensor has inherent constraints, e.g., it is limited in range, can get miscalibrated. Moreover, even for human drivers, occlusions are a common cause of accidents; for autonomous vehicles (AVs), they present an even greater challenge, as AVs lack human intuition and their perception of the environment relies entirely on sensors. When the limitations of an individual sensor suite are reached, collaboration between vehicles can compensate by aggregating information from multiple viewpoints. The current SOTA work in this area is primarily focused on collaborative perception, where sensor data from multiple vehicles is fused to construct a holistic view of the environment. This fused representation then supports downstream tasks such as object detection [593, 594], tracking [595], trajectory prediction [596], planning [597, 598], and task-invariant modules [599].
@@ 9.2 On Research Question 2
Thus, from a deployment readiness perspective, several gaps remain. Most studies lack qualitative evaluations that reveal where models fail, since metrics are typically averaged, making it difficult to extract insights about edge cases. If prediction instability occurs at certain timesteps, introducing a consistency mechanism becomes a relevant research direction. For underrepresented patterns and scenarios, more studies on cross-dataset evaluations are needed, i.e., training on one dataset and testing on another, to assess performance in out-of-distribution scenarios, which remains largely unexplored. Additionally, robustness to upstream errors (e.g., detection, tracking, localization, map quality) is insufficiently addressed in the academic community, despite its critical role in ensuring reliable real-world performance.

### v3

@@ 6.2 Anchor-based Trajectory Prediction
Anchor-based trajectory prediction methods assume that future trajectories are conditioned on planning objectives, such as goal destinations or potential paths. These methods use predefined spatial or trajectory “anchors” to guide the prediction process. Figure 12 illustrates the general pipeline for the coupled planning and prediction framework commonly found in the literature. The pipeline begins by encoding past trajectories and scene context into a context vector. This context vector is then utilized in the anchor-based decoding process to predict possible future trajectories. Anchor-based decoding can be guided by two types of anchors: goal anchors or trajectory anchors. Further we discuss each category and provide a summary of reviewed anchor-based prediction methods in Table 17.
@@ 6.2.1 Goal Anchor-Based Prediction Methods
Goal anchor-based methods [545, 85, 537, 546] focus on predicting a set of feasible goal points that serve as anchors for trajectory generation [544]. These methods predict potential goals based on the current context and agent behavior [544, 75]. Goals are often estimated based on static environmental features, including road topology, lane structures, or sidewalks. In road networks, they can be defined as lane endpoints, intersections, or turn exits, providing a structured basis for trajectory prediction. Once the goals are identified, trajectories are generated by conditioning on these anchors, often through a path-planning or motion-generation module. An advantage of goal anchor-based methods is their natural alignment with goal-driven tasks, such as reaching specific destinations.
BiTraP [188] employs a CVAE with a bi-directional LSTM decoder conditioned on multiple goal hypotheses, improving long-horizon forecasts by mitigating compounding error. SGNet [540] model estimates goals at multiple temporal scales via a stepwise goal estimator integrated into encoder-decoder architecture. PECNet [537] samples distant endpoints in latent space and uses non-local social pooling and a truncation trick to generate socially compliant long-term trajectories.
@@ 6.2.2 Trajectory Anchor-Based Prediction
Trajectory anchor-based methods rely on predefined or learned trajectory templates that represent potential future paths [3, 547, 548]. These templates serve as coarse motion representations and are selected or predicted as bases for future trajectories. The anchors can either be directly used or refined further during prediction. Common examples include straight, curved, or turning path templates. TPNet model [307] generates feasible future trajectories using polynomial curves. Inspired by the sampling-based paradigm in vehicle motion planning [549, 550], PRIME combines planning-based approaches with learning-based prediction by directly searching for reachable paths and generating potential trajectory anchors. This hybrid approach handles complex agent-map interactions while satisfying environmental and kinematic constraints [521], leveraging the strengths of both model-based and learning-based methods for optimal performance in complex scenarios. Next, to enforce constraints during trajectory generation, Li et al. [551] proposed the planning-inspired hierarchical (PiH) framework, which employs a hierarchical lateral and longitudinal decomposition to generate path and goal intentions for agents.
6.3 Summary Discussion and Future Directions
Planning-conditioned trajectory prediction integrates motion planning concepts into forecasting. Model-based methods offer interpretability and constraint satisfaction but lack adaptability, while learning-based methods handle complex behaviors yet often ignore physical limits. Anchor-based strategies guide predictions toward semantic goals but face scalability challenges in dense scenes. Future work includes combining model-based and learning-based approaches (e.g., PRIME [521]), incorporating safety constraints via differentiable control, and exploring multi-agent planning through intent modeling and graph-based reasoning.
@@ 6.3 Summary Discussion and Future Directions
An important direction in planning-conditioned trajectory prediction is the anchor generation module, which typically analyzes scene context and/or integrates spatial information to infer potential destinations or paths. However, scene context analysis can be computationally expensive and should be evaluated rigorously in terms of inference time. While map-based anchor generation is often more efficient, it must still be tested for scalability as the number of agents increases, since anchors are generated per agent. The anchor generation module is part of the prediction pipeline and should be evaluated online, rather than relying on pre-generated anchors. Thus, design of efficient and scalable anchor‑generation schemes that can leverage attention or clustering to learn compact sets and dynamically prune unlikely candidates is a direction for improvement.
@@ 7 Vision-Language Models for Prediction
Vision-Language Models (VLMs) [552, 553, 554, 555, 556, 557, 558, 559, 560, 561] are seen as promising technology that could enable vehicles and systems to deeply understand their surroundings, marking a significant advancement in the era of intelligent and explainable transportation. In the context of autonomous driving, VLMs can be roughly categorized into: Multimodal-to-Text (M2T) [562, 563, 564, 565] models, which take an image and text as the input and generate text as the output; Multimodal-to-Vision (M2V) [566, 567] models, where the input is an image and text and the output is an image; and Vision-to-Text (V2T) [568, 569] models, which take an image as the input and generate text as the output. This section mainly focuses on M2T models in the context of perception and prediction tasks. Readers are encouraged to look into the surveys [570, 571] for further details on other VLM configurations and downstream tasks. Figure 13 shows the general flow of a VLM-based prediction model. The model processes two main inputs: visual and textual data. While the image encoder processes visual inputs, e.g., single-camera images, multi-camera feeds, or LiDAR-range image, extracting spatial and object-level features; the text encoder interprets contextual information, which can include detailed instructions, scene description, specific object inquiries, navigation commands, etc. The encoded features are then passed to a cross-modal attention mechanism [572, 573, 574], which captures the relationships between visual cues and textual instructions to generate joint representations. This representation is further processed by an autonomous driving task-specific network, e.g., as for the prediction task, by a context understanding layer and decoder. The generated future trajectories consider both visual observations and semantic guidance.
@@
This section reviews VLM-based models designed for trajectory prediction and traffic scene understanding. Given the limited number of prediction models in the literature, we have included a survey of traffic scene understanding models, considering their potential integration into prediction pipelines. The section also provides a list of available language-enhanced autonomous driving datasets and discusses the challenges associated with integrating VLMs into the autonomous driving stack.
@@ 7.1 Trajectory Prediction
Table 18 summarizes reviewed VLM and LLM-based models for trajectory prediction. Yang et al. [575] proposed a two-stage evaluation framework. In the first stage, a VLM model analyzes images from six camera views and a BEV visualization with labeled agents to rank the prediction difficulty. To improve the model’s response accuracy, the VLM is provided with several examples of rankings constructed based on displacement errors from the ground truth. In the second stage, the VLM’s output is used to create a subset of training data based on the ranking, resulting in a highly representative and challenging training dataset for training conventional prediction models. Moon et al. [576] introduced VisionTrap that leverages human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. The model comprises four main components: a per-agent state encoder that generates agent-specific context features, a visual semantic encoder that encodes images combined with road segments, a text-driven guidance block that guides learning and understanding of detailed visual semantics, and a trajectory decoder. The study includes an inference-time analysis highlighting the model’s suitability for real-time deployment inference constraints.
@@
Keysan et al. [574] expanded the previously proposed CoverNet model [70] by incorporating a text-based scene representation. The prompt accepted by the model contains structured information about the target agent and its environment, constructed from polyline and Bézier curve lane representations. Peng et al. [577] introduced an explainable lane change prediction model (LC-LLM). The study reformulates the problem of lane change prediction as a language modeling task, processing driving scenario data as natural language prompts. To enhance transparency and prediction reliability, the model’s Chain-of-Thought (CoT) reasoning is refined to include explanatory elements in the prompts during the inference phase. Zheng et al. [578] leveraged LLMs to improve the global understanding of traffic context for motion prediction. The study transforms complex traffic scenarios and agents’ past trajectories into visual prompts (Transportation Context Maps) along with corresponding textual prompts. Transportation Context Maps, combined with text prompts, enable the model to generate situational insights, extracting details such as agent intentions and scenario types. The extracted information is then encoded and integrated into a conventional motion prediction pipeline.
@@ 7.2 Traffic Scene Understanding
Table 19 summarizes reviewed VLM and LLM-based models for scene understanding. Choudhary et al. [565] introduced Talk2BEV — a VLM designed to handle diverse tasks, including predicting traffic participants’ intentions, visual and spatial reasoning, and decision-making based on visual information. Romero et al. [559] argue that using VLMs to query images via natural language addresses issues of expressivity and the need for multiple single-purpose query models, but it often returns redundant and low-quality results. To improve this, they introduced Zelda, a system that adds discriminator and synonym terms to refine user queries into natural language prompts, enhancing accuracy by filtering out low-quality frames. Mo et al. [582] designed a framework for understanding driving corner cases, built upon the foundation model InternVL2 [595] and consisting of four phases. For visual perception, the framework uses DINO [596] for 2D detection and DepthAnything [597] for depth estimation. This detection information, along with the original image, is processed by a scene-aware retrieval module to generate contextually relevant information. The Driving Prompt Optimization module then uses this information to guide the model, suppressing hallucinations and focusing on critical scene aspects, ultimately producing responses for perception and driving recommendations. Zhou et al. [583] introduced the Embodied Language Model (ELM) to enhance traditional VLMs’ scene comprehension, supporting tasks such as object localization, detailed scene description, prediction, and event memorization. ELM incorporates space-aware pre-training for robust spatial localization and time-aware token selection to effectively process long-term video data and retrieve relevant past events. Finally, Tian et al. [584] proposed the DriveVLM system for planning in complex environments, with modules for scene description, scene analysis, and hierarchical planning. The scene description module generates linguistic representations of driving conditions and passes these to the scene analysis module, which examines objects’ motion states and behaviours that may influence vehicle actions.
@@ 7.4 Summary and Discussion
Another challenge is prompt engineering. It refers to the process of crafting input prompts that guide the model toward specific behaviors or outputs, especially by aligning visual and textual inputs. Providing relevant context within the prompt significantly affects the model’s response. For example, including scene-specific cues such as “heavy traffic ahead” or “intersection with pedestrians” can influence the model’s understanding and downstream reasoning. Prompts can also include instructions on output length or format, helping ensure concise and relevant generation. In some cases, few-shot prompting is used instead, where an example of the expected output is embedded within the prompt. Additionally, prompting strategies such as Chain-of-Thought (CoT) prompting guide the model to break down its reasoning into intermediate steps, enhancing both interpretability and output quality. Efficient prompt design not only improves semantic alignment but also reduces token usage, contributing to lower inference time.
Lastly, the non-deterministic nature of some VLMs [603, 604], where the same image-text input may yield different outputs across runs, presents integration challenges in safety-critical scenarios. Strategies such as adjusting the temperature parameter to control output variability, applying majority voting, or averaging across the top-$k$ outputs can help manage this variability. Striking the right balance between consistency and flexibility is essential: enforcing deterministic outputs enhances system stability, while preserving generative flexibility allows the model to better handle edge cases and out-of-distribution scenarios. A rigid deterministic setup may limit adaptability, especially in complex environments. Consequently, the choice between generative and deterministic use of VLMs must be carefully evaluated in the context of trajectory prediction. Although VLMs show promise for enhancing situational understanding and predictive reasoning, their optimal configuration within the autonomous driving stack remains an on-going research investigation.
@@ 8 Collaboration between AVs for Prediction
Despite advancements in multimodal sensing, each sensor still has fundamental limitations, such as restricted range and the risk of miscalibration. Moreover, even for human drivers, occlusions are a common cause of accidents; for autonomous vehicles (AVs), they present an even greater challenge, as AVs lack human intuition and their perception of the environment relies entirely on sensors. When the limitations of an individual sensor suite are reached, collaboration between vehicles can compensate by aggregating information from multiple viewpoints. The current SOTA work in this area is primarily focused on collaborative perception, where sensor data from multiple vehicles is fused to construct a holistic view of the environment. This fused representation then supports downstream tasks such as object detection [605, 606], tracking [607], trajectory prediction [608], planning [609, 610], and task-invariant modules [611].
@@ 9.2 On Research Question 2
Thus, from a deployment readiness perspective, several gaps remain. Most studies lack qualitative evaluations that reveal where models fail, since metrics are typically averaged, making it difficult to extract insights about edge cases. If prediction instability occurs at certain timesteps, introducing a consistency mechanism becomes a relevant research direction. For underrepresented patterns and scenarios, more studies on cross-dataset evaluations are needed, i.e., training on one dataset and testing on another, to assess performance in out-of-distribution scenarios, which remains largely unexplored.